OpenStudy (anonymous):

Write the equation of a hyperbola with vertices (3, -1) and (3, -9) and co-vertices (-6. -5) and (12, -5).

OpenStudy (anonymous):

(y+5)^2/16-(x-3)^2/81=1 (y-5)^2/16-(x+3)^2/81=1 (x-3)^2/81-(y+5)^2/16=1 (x+3)^2/81-(y-5)^2/16=1

OpenStudy (anonymous):

@jamroz

OpenStudy (anonymous):

@Cyn_A

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

\[\frac{ x^2 }{ a^2 }-\frac{ y }{ b^2 }=1\] you'll get A

OpenStudy (ghazi):

\[\frac{ y-(-5) }{ (4)^2 }-\frac{ (x-3)^2 }{ (9)^2 }=1\]